Please help me solve this linear equation by factoring: 5x-[7-(2x-1)]=3(x-5)+4(x+3)
Answer by rfer(16322)

5x-[7-(2x-1)]=3(x-5)+4(x+3)
5x-[7-2x+1]=3x-15+4x+12
5x-7+2x-1=7x-3
7x-7x=-3+6
no solution x's cancel